OpenStudy (anonymous):

What is the vertex of the graph of y = −2x2 − x + 3?

OpenStudy (anonymous):

It is a parabola b/c it has a one variable (x) raised to the second power. You need to put the equation in the standard format for a parabola to find the vertex, do you know the standard format?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

\[Vertex=(\frac{-b}{2a},f(\frac{-b}{2a}))\] If you have an equation in the form \[y=ax^2+bx+c\] Your equation is in that form. Can you tell me the values of a, b, and c?

OpenStudy (anonymous):

a= -1, b = 1, c = 3

OpenStudy (anonymous):

Your equation is \[y=-2x^2-x+3\], therefore the value of a is actually -2, the value of b is -1, and the value of c is 3, the reason is that your equation is really just: \[(-2)x^2+(-1)x+3\]

OpenStudy (anonymous):

right, because the standard form only has plus signs inbetween the terms.

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

Yeah, but that is the original equation? How do I find the vertex.

OpenStudy (anonymous):

Ok, good. To find the vertex is easy once you know the values of a,b, and c. The x coordinate of the vertex is just the formula \[\frac{-b}{2a}\], and the y co-ordinate of the vertex is just the x coordinate plugged into the original equation.

OpenStudy (anonymous):

So the x co-ordinate would be (1/(2*-2)) = -(1/4)

OpenStudy (anonymous):

Then to find the y coordinate, you just plug in -(1/4) for x into your original equation.

OpenStudy (anonymous):

So 25/8?

OpenStudy (anonymous):

I got 3.125 by plugging it into my calculator

OpenStudy (anonymous):

I need a fraction, it's 25/8

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

hopefully that works